OpenStudy (anonymous):

Identify the slope and the y-intercept of the following equation: 11x+8y=40

hartnn (hartnn):

can you isolate 'y' from that equation ? and bring it in the form of y = mx +c ?

OpenStudy (anonymous):

i thank

hartnn (hartnn):

so what you get ? y =... ?

OpenStudy (anonymous):

i have no idea im not good at matthh i listen but i still never gett it

hartnn (hartnn):

try! 11x +8y = 40 we need to isolate the 'y' so we need to bring 'x' on right side so we SUBTRACT 11x from both sides, what do u get then ?

OpenStudy (anonymous):

8y=-11x+40

hartnn (hartnn):

yes! now to isolate 'y' we just need to get rid of that 8! since its multiplied with 'y', we will need to DIVIDE both sides by 8 what do u get then ?

OpenStudy (anonymous):

11x=5

hartnn (hartnn):

how ? where did y go ? :O we have 8y = -11x + 40 dividing both sides by 8 gives us y = (-11/8)x + 40/8 which is y = (-11/8)x+ 5 now compare that with y= mx+c m will be = slope and c = y-intercept

OpenStudy (anonymous):

so the slope is (-11/8) and the y- intercept is 5

hartnn (hartnn):

absolutely correct! good :)

OpenStudy (anonymous):

can yhu help me with more

hartnn (hartnn):

ask it in a new post, so that if i am unavailable, others too can help you :)

OpenStudy (anonymous):

ok